Stock price of $42, earnings of $2.12 per share during the last twelve months, forecasted earnings of $2.84 over the following year, and average earnings growth forecast of 12.5% per year for the next five years. What is this stock's PEG, rounded to one decimal place?
Answer #1

PEG ratio using leading EPS=42/2.84*1/12.5=1.183098592

PEG ratio using trailing EPS=42/2.12*1/12.5=1.58490566